OpenStudy (anonymous):

Please help? A mountain has a rise of 8,480 feet from its base. If there is a 2 degree drop in temperature for every rise of 530 feet in altitude, how many degrees difference is there between the ground temperature and the temperature at the peak?

OpenStudy (nubeer):

530 = 1 step so 8480/ 530 this will tell u how many steps taken to reach that peak

OpenStudy (nubeer):

8480 / 530 = 16 steps. now for every step 2 degree change so 16*2 so differnce is 32 degrees

OpenStudy (anonymous):

Thank you!

OpenStudy (whpalmer4):

At the base, the temperature is \(T_g\). At some point above, the temperature will be \[T(h) = -\frac{2}{530}h+T_g\] because the temperature drops 2 degrees for every 530 feet. Plugging in the value of \(h\) at the top, we get: \[T(8480) = -\frac{2}{530}(8480) + T_g\]=-32 + T_g\] The temperature is -32 below what it was at the ground level.

OpenStudy (whpalmer4):

\[T(8480) = -32 + T_g\]

OpenStudy (anonymous):

Thanks!